a.(1-tan^2u)/(1+tan^2u)=cos(2u)
I think this is a half angle identity?verify the identity.

b. compare the graphs of y1=(1-tan^2u)/(1+tan^2u) and y2=cos(2u)

Answer by stanbon(48558)

verify the identity.
a.(1-tan^2u)/(1+tan^2u)=cos(2u)
(1-tan^2u)/sec^2u = cos^2u-sin^2u
[1/sec^2u] - [tan^2u/sec^2u] = cos^2u-sin^2u
cos^2u - sin^2u = cos^2u - sin^2u
=================================

b. compare the graphs of y1=(1-tan^2u)/(1+tan^2u) and y2=cos(2u)
---------
Can't do that on this site.
--------------------------------------
Cheers,
Stan H.